3) Do the following series diverge, absolutely converge, or conditionally converge? Write which test you use for each part.

(9 points each) (Hint: use L'Hopital's Rule for part b)

ܠܸܒܕܐܕ
Answer #1

K=2 Ratco test :- ( 1) k (akt1 -akya det an akt k K-1 kte-2 okti 1) (K- L = antl lum lem Kyo Kdo ( 1) CU len KY We 1 - HK K 7set f (x) - 2 - 1 12-11 then lem E 1 t 7D 2-1 = won log 16-11 - log (1) ( log(1) = 0. 2 um It-1) tuo Hence, The series M8 (-2